OpenStudy (anonymous):

In an examination if a student looses mark for the wrong answer and gets 1 mark for the correct answer. If he got 30 marks for solving 54 questions, then how many questions were correct?

OpenStudy (anonymous):

So, how about this: The difference in points between 54 and 30 is 24. So 54-30=24. This is how many points he was off by, this means that this is the amount of points that were canceled. So 24/2 = 12. He got 12 of those remaining 24 right, and 12 wrong. So He got 30+12 right = 42 Questions right. and 12 wrong.

OpenStudy (anonymous):

but the correct answer is 36.....

OpenStudy (dumbcow):

let r be right answers and w be wrong answers r+w = 54 r-w = 30 adding equations: 2r = 84 r = 42 42 right , 12 wrong

OpenStudy (dumbcow):

is the question posted correctly? loses 1 mark for wrong answer and gains 1 mark for right answer ?

OpenStudy (dumbcow):

the only way you end up with 30 marks with 36 right answers means he loses (1/3) mark for each wrong answer